write the first four terms of each sequence
whose general term is given.
an=(-1/3)^n

Answer by stanbon(75887)

write the first four terms of each sequence
whose general term is given.
an=(-1/3)^n
----
a(1) = (-1/3)^1 = -1/3
a(2) = (-1/3)^2 = 1/9
a(3) = (-1/3)^3 = -1/27
a(4) = (-1/3)^4 = 1/81
a(5) = (-1/3)^5 = -1/243
----------------------------
Cheers,
Stan H.
-------------------